450 tickets were sold for a game for a total of $562.50. If adult tickets sold for $1.50 and children's tickets sold for $1.00, how many of each kind of ticket were sold?
Answer by Cromlix(4381)

Hi,
Make adult = x
Make child = y
x + y = 450......Equation (1)
1.5x + y = 562.50....Equation (2)
Take Eq(1) from Eq(2)
0.5x = 112.50
x = 225
substitute x = 225 into Eq(1)
x + y = 450
225 + y = 450
y = 225
225 Adult's tickets were sold
225 children's tickets were sold
Hope this helps :-)